OpenStudy (milo123):

D? The legs of a right triangle are 3 units and 2 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. 1.00 2.24 3.61 5.00

OpenStudy (milo123):

I think C or D?

OpenStudy (3mar):

 Pythagorean Theorem

OpenStudy (milo123):

D

OpenStudy (3mar):

\[\sqrt{2^2+3^3}=\sqrt{13}=3.6055513\]

OpenStudy (3mar):

Did you get it?

OpenStudy (milo123):

wait so it's c?

OpenStudy (3mar):

Great! Yes it is C (3.6)

OpenStudy (3mar):

soryy 3.61